# Domain Adaptation for Contrastive Audio-Language Models

### v1

@@ 1 Introduction
The human prompt engineering improves the ALM's performance on the specific domain. For example, manual prompt engineering [6] from ``{class}" to ``this is a sound of {class}" leads to a relative 2% improvement in performance on audio event datasets. The text prompts are designed for each domain using an ad-hoc experimental process. This process generally requires knowledge of the ALM's training data which may or may not be available to the public. Existing approaches improve domain performance by Few-Shot Learning (FSL). This involves learning prompts [7] or adapters [8] using supervised training on annotated target data. The annotation process is expensive and may not be feasible for deployed models. FSL methods also lead to a loss in generalization of base ALM and poor out-of-distribution performance [9, 8].
@@
In this work, we propose a domain adaptation method for Contrastive ALMs that is performed at test-time using unlabeled audio. Our method involves learning a domain vector, which helps the model adjust to new audio domains by enforcing consistent prediction across augmented views of audio. This process is divided into three stages: Augment: The test audio is subject to various augmentations to simulate different listening conditions. Alongside this, a domain vector is incorporated into the model's text processing component; b) Combine: The model makes predictions based on the augmented audio and the modified text embeddings. These predictions are then averaged to produce a more stable and generalized output, and c) Optimize: Finally, the averaged prediction probabilities are used to compute self-entropy to fine-tune the domain vector, ensuring that the model's predictions are adapted for the domain. With only one unlabelled audio, we see an average 3.2% (max 8.4%) zero-shot performance improvement across 12 downstream tasks, while retaining the generalization property of ALMs.
@@ 2.1 Contrastive Audio-Language Models
This section provides a generic overview of Contrastive Audio-Language models in literature. We will commonly refer to them as ``model". The model has a two-tower structure consisting of audio and text branches. The two branches are pretrained using contrastive learning [15]. After pretraining, the model can be used for zero-shot classification. To illustrate, the user provides an audio file and the $N$ classes to classify the audio into. The audio file is converted into a mel-spectrogram ($x\in\mathbb{R}^{T\times F}$) and passed to an audio encoder to produce audio embeddings $v\in\mathbb{R}^{1\times d}$. On the text side, $N$ classes are tokenized and embeddings are obtained for each token. Then the token embeddings are passed through a text encoder to produce text embeddings $u\in\mathbb{R}^{d\times N}$. Both embeddings are projected into a common embedding space using two independent projection layers. After projection, the dot product is computed between the two embeddings. Generally, an activation function like Softmax or sigmoid is applied to obtain a valid probability distribution across $N$ classes of interest.
@@ 2.2 Human-crafted text-prompts
The contrastive models like CLAP [6] append prompts to classes for better zero-shot performance. The prompts are hand-crafted for the domain the model is being tested on. For example, a prompt of "this emotion is " is used for Speech Emotion Recognition while the prompt of "this is the sound of " is used for Sound Event Classification. The choice of text prompt for each task affects performance [6, 16]. From the prompt experiment, we hypothesize that: though audio-language models have an adequate understanding of audio concepts, we don't have methods to elicit this information from the models. The manual method of hand-crafting prompts is cumbersome and requires domain knowledge. Moreover, the handcrafted prompts don't take into account intra-task variance resulting from domain changes. For example, zero-shot inference for the task of sound event classification can provide different performances when evaluated in an urban city, a domestic household versus in a rainforest. This leads us to believe an automatic unsupervised adaptation method for zero-shot inference of audio-language can help improve performance.
@@ 3.1 Augment
Time Reorder (TR): Audio-language models compress all audio information into a single embedding which leads to loss of temporal information [18]. Therefore, we explicitly add a time-reorder augmentation.
@@
Table 1: Zero-shot and Test-time adaptation performance on 12 downstream tasks. The adaptation method uses one unlabelled audio example and five unlabelled audio examples at test time. Higher is better for all numbers. The metric for all the tasks is Accuracy. Average is the average score of all downstream tasks
@@ 5.2 Effect of Augmentations
Zero-Shot 25 Augment 50 Augment
One example 62.93 63.51 64.95
Five example 65.92 63.77 65.91
@@ 5.4 Limitations
Our adaptation method improves zero-shot performance by using one unlabelled audio at test-time. Its practical applications include on-device updates for audio-language model model to adapt it to the user’s audio and media. In this setup, access to labeled annotations is not possible but the model does have access to test audio used for inference. However, compared to zero-shot inference, the method incurs a higher computation cost due of back-propagation to update the domain vector and processing multiple augmented views. Depending on the downstream application and the on-device compute, the parameter update may not be feasible, limiting the applicability of this method.
@@ 6 Conclusion
We propose a domain adaptation method for ALMs. The method uses unlabelled audio at test-time to enforce consistent prediction across augmented views of audio by optimizing a domain vector. It consists of three parts: Augment, Combine, and Optimise. Augment: The test audio is subject to various augmentations to simulate different listening conditions. Alongside this, a domain vector is incorporated into the model's text processing component; b) Combine: The model makes predictions based on the augmented audio and the modified text embeddings. These predictions are then averaged to produce a more stable and generalized output, and c) Optimize: Finally, the averaged prediction probabilities are used to compute self-entropy to fine-tune the domain vector, ensuring that the model's predictions are adapted for the domain. Adaption with one unlabelled audio provides an average 3.2% (max 8.4%) zero-shot improvement and with 5 audios an average improvement of 4.7% (max 9.4%). After adaptation, the ALM still retains its zero-shot generalization.

### v2

@@ 1 Introduction
The human prompt engineering improves the ALM’s performance on the specific domain. For example, manual prompt engineering [6] from {class}" to “this is a sound of {class}" leads to a relative 2% improvement in performance on audio event datasets. The text prompts are designed for each domain using an ad-hoc experimental process. This process generally requires knowledge of the ALM’s training data which may or may not be available to the public. Existing approaches improve domain performance by Few-Shot Learning (FSL). This involves learning prompts [7] or adapters [8] using supervised training on annotated target data. The annotation process is expensive and may not be feasible for deployed models. FSL methods also lead to a loss in generalization of base ALM and poor out-of-distribution performance [9, 8].
@@
In this work, we propose a domain adaptation method for Contrastive ALMs that is performed at test-time using unlabeled audio. Our method involves learning a domain vector, which helps the model adjust to new audio domains by enforcing consistent prediction across augmented views of audio. This process is divided into three stages: Augment: The test audio is subject to various augmentations to simulate different listening conditions. Alongside this, a domain vector is incorporated into the model’s text processing component; b) Combine: The model makes predictions based on the augmented audio and the modified text embeddings. These predictions are then averaged to produce a more stable and generalized output, and c) Optimize: Finally, the averaged prediction probabilities are used to compute self-entropy to fine-tune the domain vector, ensuring that the model’s predictions are adapted for the domain. With only one unlabelled audio, we see an average 3.2% (max 8.4%) zero-shot performance improvement across 12 downstream tasks, while retaining the generalization property of ALMs.
@@ 2.1 Contrastive Audio-Language Models
This section provides a generic overview of Contrastive Audio-Language models in literature. We will commonly refer to them as “model". The model has a two-tower structure consisting of audio and text branches. The two branches are pretrained using contrastive learning [15]. After pretraining, the model can be used for zero-shot classification. To illustrate, the user provides an audio file and the $N$ classes to classify the audio into. The audio file is converted into a mel-spectrogram ($x\in\mathbb{R}^{T\times F}$) and passed to an audio encoder to produce audio embeddings $v\in\mathbb{R}^{1\times d}$. On the text side, $N$ classes are tokenized and embeddings are obtained for each token. Then the token embeddings are passed through a text encoder to produce text embeddings $u\in\mathbb{R}^{d\times N}$. Both embeddings are projected into a common embedding space using two independent projection layers. After projection, the dot product is computed between the two embeddings. Generally, an activation function like Softmax or sigmoid is applied to obtain a valid probability distribution across $N$ classes of interest.
@@ 2.2 Human-crafted text-prompts
The contrastive models like CLAP [6] append prompts to classes for better zero-shot performance. The prompts are hand-crafted for the domain the model is being tested on. For example, a prompt of "this emotion is " is used for Speech Emotion Recognition while the prompt of "this is the sound of " is used for Sound Event Classification. The choice of text prompt for each task affects performance [6, 16]. From the prompt experiment, we hypothesize that: though audio-language models have an adequate understanding of audio concepts, we don’t have methods to elicit this information from the models. The manual method of hand-crafting prompts is cumbersome and requires domain knowledge. Moreover, the handcrafted prompts don’t take into account intra-task variance resulting from domain changes. For example, zero-shot inference for the task of sound event classification can provide different performances when evaluated in an urban city, a domestic household versus in a rainforest. This leads us to believe an automatic unsupervised adaptation method for zero-shot inference of audio-language can help improve performance.
@@ 3.1 Augment
Time Reorder (TR): Audio-language models compress all audio information into a single embedding which leads to loss of temporal information [18]. Therefore, we explicitly add a time-reorder augmentation (details in appendix).
@@
Table 1: Zero-shot and Test-time adaptation performance on 12 downstream tasks. The adaptation method uses one randomly chosen unlabelled audio example and five randomly chosen unlabelled audio examples at test time. The Table reports the average numbers across 5 runs. The metric is accuracy and higher is better for all tasks. Average $\uparrow$ is the average performance across 12 tasks
@@ 5.2 Effect of Augmentations
One test example Five test example
25 Augmentations 63.51 63.77
50 Augmentations 64.94 65.92
@@ 5.4 Limitations
Our adaptation method improves zero-shot performance by using one unlabelled audio at test-time. Its practical applications include on-device updates for audio-language model model to adapt it to the user’s audio and media. In this setup, access to labeled annotations is not possible but the model does have access to test audio used for inference. However, compared to zero-shot inference, the method incurs a higher computational cost because of backpropagation used to update the domain vector and the processing of multiple augmented views. Depending on the downstream application and the on-device compute, the parameter update may not be feasible, limiting the applicability of this method.
@@ 6 Conclusion
We propose a domain adaptation method for ALMs. The method uses unlabelled audio at test-time to enforce consistent prediction across augmented views of audio by optimizing a domain vector. It consists of three parts: Augment, Combine, and Optimise. Augment: The test audio is subject to various augmentations to simulate different listening conditions. Alongside this, a domain vector is incorporated into the model’s text processing component; b) Combine: The model makes predictions based on the augmented audio and the modified text embeddings. These predictions are then averaged to produce a more stable and generalized output, and c) Optimize: Finally, the averaged prediction probabilities are used to compute self-entropy to fine-tune the domain vector, ensuring that the model’s predictions are adapted for the domain. Adaption with one unlabelled audio provides an average 3.2% (max 8.4%) zero-shot improvement and with 5 audios an average improvement of 4.7% (max 9.4%). After adaptation, the ALM still retains its zero-shot generalization.